OpenStudy (anonymous):

how many moles are in 500ml of he gas at stp

OpenStudy (anonymous):

.0223 moles

OpenStudy (abb0t):

@rman3 can you explain how you got that answer instead.

OpenStudy (anonymous):

Use PV=nrt, the ideal gas law. P=Pressure v=volume in liters n=moles r=gas constant(.0821) = t=temp in K. n=PV/RT n=(1)(.5)/(.0821)(273) n=.0223

OpenStudy (anonymous):

Many ways to do a problem man

OpenStudy (abb0t):

\(\huge NO\)

OpenStudy (anonymous):

YEAH

OpenStudy (anonymous):

since the gas is at STP, you can use this formula..... \[n = \frac{ V }{ V_M }\] n=no. of moles V=Volume V(M)=Molar volume (volume occupied by one mole of a gas at STP ) and this equals 24dm^3 = 24L